OpenStudy (anonymous):

Which point below would you use with the point (-2, -5) to get a line with a slope of -1? a) (-6, 1) b) (-1, -6) c) (1, 4) d) (4, 1)

hero (hero):

Use the slope formula to set this up as follows: \[m = \frac{y_2 - y_1}{x_2-x_1}\]\[-1 = \frac{y + 2}{x+5}\]Then input one of the points until it satisfies the equation

hero (hero):

After doing a little manipulation, I get -7 = x + y So whichever point adds up to -7 is correct

hero (hero):

looks like b

OpenStudy (anonymous):

I'm looking over it to understand it. Thank you :]

hero (hero):

Good luck with that :D